# Verifying LLM Inference to Detect Model Weight Exfiltration

### v2

@@ 5.1 The Cost of Verification versus Generation
During verification, a single token mismatch between two generations makes their subsequent probability distributions divergent, as they are conditioned on different contexts (e.g., all continuations of “the cat sat on” vs. “the cat sat beside” will diverge). Instead of continuing togenerating text with fixed seeds—which would diverge immediately—we exploit the fact that LLMs expose logits for all next-token predictions during the prefill stage. A single prefill over the full input–output sequence yields conditional probabilities for every position, enabling token-level suspiciousness scores in a single forward pass. Verification therefore involves only prefill, whereas normal inference requires both prefill and autoregressive decode. In short-context regimes, decoding achieves substantially lower efficiency than $3–5\times$ lower hardware utilization is typical for short-context decoding than prefill, which implies a $\penalty 10000\ 3–5\times$ throughput gap. This gap increases in long context regimes, although the exact difference will depend on attention optimizations and implementation details [erdil2025inferenceeconomicslanguagemodels].
@@ 5.2 Inverse Probability Transform Scoring Function Token-IPT-DiFR
In Section 4.5 we introduce the notion of a classification using Fixed-Seed Sampling Likelihood. In this section we introduce a technique for approximating the fixed-seed sampling likelihood (FSSL) for the inverse probability transform sampling technique (IPT). We investigate a method Token-IPT-DiFR, which is concurrently introduced in parallel work [tokendifr]. The core assumption is that across repeated model evaluations, the per-token probability vectors $\vec{p}_{\theta}(x)$ may vary due to small numerical or stochastic perturbations, but such variation is relatively small.
Intuitively, if the true sampling process (with fixed seed $\sigma$) selects a random value $u\in[0,1]$ that falls within the probability interval of one token, nearby tokens—whose cumulative probability mass lies close to uu—remain plausible under valid non-determinism. Hence, instead of measuring the full empirical fixed-seed posterior (which would require many re-samplings per context), we approximate the likelihood of plausible tokens with a Gaussian centered at the seed sample $u$.
Formally, we compute a verification score for the observed token $t^{\star}$ by integrating a Gaussian density (mean $u$, width $\sigma$) over the token’s probability interval $[a,b]$ in the cumulative distribution $\bar{F}$ of $\vec{p}_{\theta}(x)$. This integral estimates the probability that a perturbed distribution—drawn from $\mathcal{P}_{\theta}(x)$—would produce $t^{\star}$ under the same seed. The log of this mass gives a continuous, differentiable proxy for the fixed-seed posterior probability. The Token-IPT-DiFR is defined in Algorithm 3.
Figure 7: Diagram of the Token-IPT-DiFR method
@@ 5.3 Gumbel-Max Scoring Function Token-DiFR
We also introduce the Token-DiFRto estimate FSSL-GM, which is defined for a given noise distribution over token probabilities, which by default we assume to be Gaussian $\mathcal{N}$, also introduced in concurrent work [tokendifr].
Given a sampled token $t^{*}$, and knowledge of the inference server’s sampling seed $\sigma$, the verifier regenerates the same Gumbel noise the inference server generated $\vec{g}_{\sigma}\sim\mathrm{Gumbel}(0,1)^{V}$. The verification server generates the probability distribution $\hat{\ell}$. Then, we compute the probability that token $t^{*}$ surpassed that token as the sampled token; that is, computing the probability $\Pr[\vec{g}_{\sigma}[i]+\hat{\ell}[i]+\mathcal{N}>\vec{g}_{\sigma}[t^{*}]+\hat{\ell}[t^{*}]+\mathcal{N}]$ for token $i$. The full method and derivations are provided in Appendix B.2, and Algorithm 4.
Figure 8: Diagram of what scalar the Token-DiFR function aims to capture, in order to estimate the likelihood of a particular token being sampled, given a specific noise distribution across token probabilities.
@@ 5.4 How to Pool Across Multiple Tokens
In practice, computing a per-token suspiciousness score is only the first step—the harder problem is determining how to pool these scores across larger units of activity. Apriori, as a defender we do not know how an adversary will distribute an exfiltration attempt (e.g., they can attack across all tokens from a single user, across the first token of many users, or across all outputs generated on a particular GPU); thus there is no universally correct aggregation rule. Simple pooling strategies such as taking the 90th percentile or the mean correspond to different implicit threat models. The security team must define the granularity at which suspicious tokens are acted upon (e.g. user-level, request-level, or hardware-level) and the time horizon over which evidence accumulates. While our method provides calibrated token-level scores, the pooling and escalation policy is necessarily system-dependent and must be tailored to the deployment’s trust boundaries and adversary model. In this work we are purposefully agnostic to the adversary and thus do not comment on the right method to pool scores across tokens.
@@ 6.1 Problems with FSSL Estimators and Logit-Rank as a Solution
$\textsc{safe}:\ \mathrm{FSSL}_{t}(i)\!\geq\!\tau\quad\textsc{suspicious}:\ \mathrm{FSSL}_{t}(i)\!<\!\tau\ \wedge\ \mathrm{rank}_{t}(i)\!\leq\!R\quad\textsc{dangerous}:\ \mathrm{rank}_{t}(i)\!>\!R$
@@
$\mathrm{ExfiltratableInformation}(t)=\begin{cases}\log_{2}|A_{t}|,&\text{if }C(t)=\textsc{safe},\\[6.0pt] \log_{2}R,&\text{if }C(t)=\textsc{suspicious},\\[6.0pt] \log_{2}V,&\text{if }C(t)=\textsc{dangerous}.\end{cases}$
@@
We sample 500 prompts from LMSYS-Chat-1M, and for each prompt, we generate continuations of max length 500 tokens. We evaluate Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, Qwen3-1.5-MoE-A.2-7B, and Qwen3-30B-A3B, served via vLLM. [^7] We evaluate on vLLM as this is the most popular open-source inference provider, and thus aim toe stimate FSSL-GM scores, rather than FSSL-IPT, as vLLM uses the Gumbel-Max trick for sampling. For all our experiments we use Token-DiFR. We use temperature $T=1.$, top-$k=50$, top-$p=0.95$, and a fixed RNG seed for generation. The verifier shares the tokenizer and chat template with the generator. We sweep over multiple $\sigma$ values (from 0.001 to 1.) for Token-DiFR hyperparameters.
We focus on verifying inference conducted via the standard auto-regressive decoding process. While speculative decoding is increasingly common in practice, our verification pipeline is designed around the traditional decoding setup; we discuss extensions to speculative decoding in Appendix E.1.
@@
In Section H.1, we demonstrate an ablation of these findings and show a robust improvement when introducing logit rank, where if rank =None then all tokens below an FSSL threshold $\tau$ are classified as dangerous. More information on the FSSL-GM scores for different tokens are presented in Appendix G.
@@ 7 Discussion
One of core challenges of this proposal is that if an adversary is able to corrupt an inference provider, then meaningful steps must be taken to protect the verification server from the same threat. There are multiple possible solutions to this, depending on the specific threat model an inference provider may have. In the most extreme setting, an inference provider may create a secondary inference stack for verifying inference that is kept in parallel with the main inference stack, it may even run in a Trusted Execution Environment to verify integrity of the computation. In the event that this is too costly for an inference provider to maintain, an earlier code checkpoint from a point prior to when an adversary was known to be in the system, can be used as a reference implementation. Further, the defending the verification server benefits from not having the same performance requirements of the inference stack — depending on availability constraints, the verification server could even exist in a largely air-gapped format with minimal connection to the inference stack, except at limited times across a secure channel. The fact that verification is much more efficient than generation, and that randomly sampling messages to verify can produce high confidence probabilistic guarantees (Section C) suggests that the amount of systems-level overhead required to implement a verification server may not be a blocker, if steganography is deemed a meaningful threat.
This system cannot standalone as a method for defending against against model weight exfiltration; in Appendix J, we describe how this scheme fits into a larger outline of defenses to protect model weights. At the very least, this scheme only detects steganographic messages, and does not actually prevent model weight exfiltration. There are a series of responses that are possible, in response to suspicious tokens; for example, tokens deemed suspicious can be logged and aggregated to look for potentially malicious patterns by an adversary. Alternatively, outbound traffic can be gated by this verification server, and when failing verification, inference is regenerated and tried again. It is possible to rate limit users (or GPUs) proportional to the number of tokens that they produce that are deemed suspicious. We give a more complete discussion of policy recommendations of how this scheme fits into a larger scheme of in Appendix F.
@@ Appendix A Appendix
Appendix B Algorithms
@@
B.1 Inverse Probability Transform (IPT) Fixed-Seed Sampling Likelihood (FSSL) Estimators
@@
Algorithm 3 Token-IPT-DiFR (from model $\theta$ and context)
@@
1:model parameters $\theta$;token probabilities $p\in\Delta^{|V|}$; observed token $t^{\star}\in\{1,\dots,|V|\}$; seed $s$; width $\sigma>0$; small $\varepsilon>0$
@@
[…]
10:return $S$
B.2 Gumbel-Max FSSLEstimators (FSSL-GM)

### v3

@@ 6.1 The Cost of Verification versus Generation
[…]
@@ 6.2 Inverse Probability Transform Scoring Function Token-IPT-DiFR
In Section 5.2 we introduce the notion of a classification using Fixed-Seed Sampling Likelihood. In this section we introduce a technique for approximating the fixed-seed sampling likelihood (FSSL) for the inverse probability transform sampling technique (IPT). We investigate a method Token-IPT-DiFR, which is concurrently introduced in parallel work [KRR+25]. The core assumption is that across repeated model evaluations, the per-token probability vectors $\vec{p}_{\theta}(x)$ may vary due to small numerical or stochastic perturbations, but such variation is relatively small.
Intuitively, if the true sampling process (with fixed seed $\sigma$) selects a random value $u\in[0,1]$ that falls within the probability interval of one token, nearby tokens, whose cumulative probability mass lies close to $u$, remain plausible under valid non-determinism. Hence, instead of measuring the full empirical fixed-seed posterior (which would require many re-samplings per context), we approximate the likelihood of plausible tokens with a Gaussian centered at the seed sample $u$.
Formally, we compute a verification score for the observed token $t^{\star}$ by integrating a Gaussian density (mean $u$, width $\sigma$) over the token’s probability interval $[a,b]$ in the cumulative distribution $\bar{F}$ of $\vec{p}_{\theta}(x)$. This integral estimates the probability that a perturbed distribution, drawn from $\mathcal{P}_{\theta}(x)$, would produce $t^{\star}$ under the same seed. The log of this mass gives a continuous, differentiable proxy for the fixed-seed posterior probability. The Token-IPT-DiFR is defined in Algorithm 3, and illustrated in Figure 10.
@@ 6.3 Gumbel-Max Scoring Function Token-DiFR
We also introduce the Token-DiFRto estimate FSSL-GM, which is defined for a given noise distribution over token probabilities, which by default we assume to be Gaussian $\mathcal{N}$, also introduced in concurrent work [KRR+25].
Given a sampled token $t^{*}$, and knowledge of the inference server’s sampling seed $\sigma$, the verifier regenerates the same Gumbel noise the inference server generated $\vec{g}_{\sigma}\sim\mathrm{Gumbel}(0,1)^{V}$. The verification server generates the probability distribution $\hat{\ell}$. Then, we compute the probability that token $t^{*}$ surpassed that token as the sampled token; that is, computing the probability $\Pr[\vec{g}_{\sigma}[i]+\hat{\ell}[i]+\mathcal{N}>\vec{g}_{\sigma}[t^{*}]+\hat{\ell}[t^{*}]+\mathcal{N}]$ for token $i$. The full method and derivations are provided in Appendix C.2, and Algorithm 4.
Figure 6: Diagram of the Token-DiFR function. The Token-DiFR estimates the likelihood of a particular token being sampled given a specific noise distribution across token probabilities.
@@ 6.4 How to Pool Across Multiple Tokens
In practice, computing a per-token suspiciousness score is only the first step; the harder problem is determining how to pool these scores across larger units of activity. Apriori, as a defender we do not know how an adversary will distribute an exfiltration attempt (e.g., they can attack across all tokens from a single user, across the first token of many users, or across all outputs generated on a particular GPU); thus there is no universally correct aggregation rule. Simple pooling strategies such as taking the 90th percentile or the mean correspond to different implicit threat models. The security team must define the granularity at which suspicious tokens are acted upon (e.g. user-level, request-level, or hardware-level) and the time horizon over which evidence accumulates. While our method provides calibrated token-level scores, the pooling and escalation policy is necessarily system-dependent and must be tailored to the deployment’s trust boundaries and adversary model. In this work we are purposefully agnostic to the adversary and thus do not comment on the right method to pool scores across tokens.
6.5 Message Sampling Rates
To understand the efficiency of our verification scheme, we analyze what percentage of traffic must be sampled to detect malicious messages with high confidence. Figure 7 illustrates the key statistical property of our approach: the sampling percentage required depends primarily on the proportion of malicious messages in the traffic. In Figure 7 we show the percentage of messages that need to be sampled in order to sample at least 1000 malicious messages, assuming the malicious messages are IID. Notably, even for very slow exfiltration attempts (e.g., 1% malicious traffic), the sampling percentage remains practical, even for high confidence. See Appendix D for the full derivation.
@@ 7.1 Problems with FSSL Estimators and Logit-Rank as a Solution
$\textsc{safe}:\ \mathrm{FSSL}_{t}(i)\!\geq\!\tau$
$\textsc{suspicious}:\ \mathrm{FSSL}_{t}(i)\!<\!\tau\ \wedge\ \mathrm{rank}_{t}(i)\!\leq\!R$
$\textsc{dangerous}:\ \mathrm{FSSL}_{t}(i)\!<\!\tau\ \wedge\ \mathrm{rank}_{t}(i)\!>\!R$
@@
$\mathrm{ExfiltratableInformation}(t)=\begin{cases}\log_{2}|A_{t}|,&\text{if }C(t)=\textsc{safe},\\[6.0pt] \log_{2}R,&\text{if }C(t)=\textsc{suspicious},\\[6.0pt] \log_{2}|\mathcal{T}|,&\text{if }C(t)=\textsc{dangerous}.\end{cases}$
@@
We sample 500 prompts from LMSYS-Chat-1M [ZCS+24], and for each prompt, we generate continuations of max length 500 tokens. We evaluate Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct [AI@24], Qwen3-1.5-MoE-A.2-7B, and Qwen3-30B-A3B [YLY+25], served via vLLM. [^6] We evaluate on vLLM as this is the most popular open-source inference provider, and thus aim to estimate FSSL-GM scores, rather than FSSL-IPT, as vLLM uses the Gumbel-Max trick for sampling. For all our experiments we use Token-DiFR. We use temperature $T=1.$, top-$k=50$, top-$p=0.95$, and a fixed RNG seed for generation. The verifier shares the tokenizer and chat template with the generator. We sweep over multiple $\sigma$ values (from 0.001 to 1.) for Token-DiFR hyperparameters.
We focus on verifying inference conducted via the standard auto-regressive decoding process. While speculative decoding is increasingly common in practice, our verification pipeline is designed around the traditional decoding setup; we discuss extensions to speculative decoding in Appendix F.1.
@@
In Section I.1, we demonstrate an ablation of these findings and show a robust improvement when introducing logit rank, where if rank =None then all tokens below an FSSL threshold $\tau$ are classified as dangerous. More information on the FSSL-GM scores for different tokens are presented in Appendix H.
@@ 8 Discussion
One of core challenges of this proposal is that if an adversary is able to corrupt an inference provider, then meaningful steps must be taken to protect the verification server from the same threat. There are multiple possible solutions to this, depending on the specific threat model an inference provider may have. In the most extreme setting, an inference provider may create a secondary inference stack for verifying inference that is kept in parallel with the main inference stack, it may even run in a Trusted Execution Environment to verify integrity of the computation. In the event that this is too costly for an inference provider to maintain, an earlier code checkpoint from a point prior to when an adversary was known to be in the system, can be used as a reference implementation. Further, the defending the verification server benefits from not having the same performance requirements of the inference stack ; depending on availability constraints, the verification server could even exist in a largely air-gapped format with minimal connection to the inference stack, except at limited times across a secure channel. The fact that verification is much more efficient than generation, and that randomly sampling messages to verify can produce high confidence probabilistic guarantees (Section D) suggests that the amount of systems-level overhead required to implement a verification server may not be a blocker, if steganography is deemed a meaningful threat.
This system cannot standalone as a method for defending against against model weight exfiltration; in Appendix L, we describe how this scheme fits into a larger outline of defenses to protect model weights. It’s important to note that this scheme does not prevent model weight exfiltration and rather only detects steganographic messages. There are a set of possible security responses in response to suspicious tokens; for example, tokens deemed suspicious can be logged and aggregated to look for potentially malicious patterns by an adversary. Alternatively, outbound traffic can be gated by this verification server, and when failing verification, inference is regenerated and tried again. It is possible to rate limit users (or GPUs) proportional to the number of tokens that they produce that are deemed suspicious. We give a more complete discussion of policy recommendations of how this scheme fits into a larger scheme of in Appendix G.
@@ Appendix A Appendix
Appendix B Steganography Syntax and Relevant Results
In this section, we review some relevant notation, concepts, and terms that are used to understand steganographic protocols. We also reprove some well known results that support our own theorems in the main body of the paper.
Definition B.1 (Steganographic Scheme).
Let $\ell_{\mathrm{msg}},\ell_{\mathrm{cov}}\in\mathbb{N}$ be integers and $\mathcal{T}$ be a set of tokens (a.k.a. vocabulary or alphabet). Let $\mathcal{D}$ be the passive channel distribution with support $X\subseteq\mathcal{T}^{\leq\ell_{\mathrm{cov}}}$. A steganography scheme for $\mathcal{D}$ is a pair of stateful (possibly randomized) algorithms $\mathsf{Steg}=(\mathsf{Enc},\mathsf{Dec})$ with the following syntax:
  • $\mathsf{Enc}(K,\mathsf{st},m)\to c$, takes a key $K\in\{0,1\}^{\lambda}$, a state $\mathsf{st}$, a message $m\in\{0,1\}^{\leq\ell_{\mathrm{msg}}}$ and outputs a covertext $c\in\mathcal{T}^{\leq\ell_{\mathrm{cov}}}$.
  • $\mathsf{Dec}(K,\mathsf{st},c)\to m$, takes a key $K\in\{0,1\}^{\lambda}$, a state $\mathsf{st}$, a covertext $c\in\mathcal{T}^{\leq\ell_{\mathrm{cov}}}$ and outputs a message $m\in\{0,1\}^{\leq\ell_{\mathrm{msg}}}$.
For a steganography scheme to be useful, it must be correct, meaning that for every $\mathsf{st}$ and $m$, $\Pr_{K}[\mathsf{Dec}(K,\mathsf{st},\mathsf{Enc}(K,\mathsf{st},m))=m]=1-\mathrm{negl}(\lambda).$ When, this probability is $1$, we say the scheme is perfectly correct. It must also be secure, meaning that for any $\mathsf{st}$ and $m$, $D_{\mathrm{KL}}(\mathcal{D}_{\mathsf{Steg}}\|\mathcal{D})=\mathrm{negl}(\lambda),$ where $\mathcal{D}_{\mathsf{Steg}}$ is the covertext distribution generated as $K\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\{0,1\}^{\lambda}$ and $c\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\mathsf{Enc}(K,\mathsf{st},m)$.
The principal measure of efficiency that we are interested in is the rate of a scheme $\mathsf{Steg}$. This is defined as the message length $\ell_{\mathrm{msg}}$ divided by the average covertext length, which we calculate as $\mathbb{E}_{c\sim\mathcal{D}}[|c|]$, since a secure scheme cannot differ significantly from this expectation. [^7] This is assuming that one requires correctness for a message distribution which is chosen uniformly at random and is therefore incompressible. Therefore, we write $R=\ell_{\mathrm{msg}}/\mathbb{E}_{c\sim\mathcal{D}}[|c|]$.
Note that there are many other variants of steganography, including stateless, public-key, robust, etc. [AP02, HLV02, VH04]. However, in our context, we wish to give a potential exfiltrator as much theoretical power as possible. So we assume that both endpoints are stateful and can have an arbitrarily long secret key in common.
B.0.1 Relevant Results in Steganography
Theorem B.2 (Steganographic limitations).
Let $\mathsf{Steg}$ be a steganography scheme for $\mathcal{D}$ and $\mathbf{M}$ be a random variable with $H(\mathbf{M})=\ell$. Let $\mathcal{D}_{\mathsf{Steg}}$ be the covertext distribution generated as $K\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\{0,1\}^{\lambda}$ and $c\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\mathsf{Enc}(K,\mathsf{st},\mathbf{M})$. Then, $\mathsf{Steg}$ cannot be perfectly correct if $\ell>\lceil H(\mathcal{D}_{\mathsf{Steg}})\rceil.$
Proof.
This theorem is an immediate consequence of Shannon’s source coding theorem [SHA48, So59]. This theorem states: (i) that no random variable $X$ can be encoded with fewer bits than $H(X)$ in expectation and (ii) that there exists an encoding of $X$ with expected length at most $\lceil H(X)\rceil$. Therefore, if $\lceil H(\mathcal{D}_{\mathsf{Steg}})\rceil<H(\mathbf{M})$ and $\mathsf{Steg}$ is correct with probability $1$, then one could encode $\mathbf{M}$ as the covertext, using a random key $K$ and random state $\mathsf{st}$, $\mathbf{C}\leftarrow\mathsf{Enc}(K,\mathsf{st},\mathbf{M})$ and decoded with $\mathsf{Dec}(K,\mathsf{st},\mathbf{C})$. Then, based on the source coding theorem, $\mathbf{C}\sim\mathcal{D}_{\mathsf{Steg}}$ (and therefore $\mathbf{M}$) can be encoded with $\lceil H(\mathcal{D}_{\mathsf{Steg}})\rceil<H(\mathbf{M})$ bits, which is a direct contradiction. ∎
In contrast to the above theorem and to illustrate the difficulty of preventing covert exfiltration, we recall a well-known theorem from [HLV02] stated in Theorem B.3. This shows that assuming classical cryptography exists at all, i.e., one-way functions exist, then one can construct a steganographic scheme that is near-optimal. The following theorem shows that there is some steganographic scheme which fools any efficient policy under minimal assumptions. Here and the rest of the paper, a policy is a function $\mathcal{W}:X(\mathcal{D})\to\{0,1\}$, which aims to output $1$ on covertexts generated steganographically but not on samples from the honest distribution. We say that $\mathcal{W}$ has a false positive rate $\alpha$ over $\mathcal{D}$ if $\Pr_{x\sim\mathcal{D}}[\mathcal{W}(x)=1]=\alpha$.
Theorem B.3 (Due to [HLV02]).
Let $\mathcal{W}$ be an efficiently-computable policy with a false positive rate $\alpha$ over $\mathcal{D}$. Then, assuming a $\delta$-secure pseudorandom function $\mathsf{PRF}$ exists, there is an inefficient steganography scheme $\mathsf{Steg}$ for $\mathcal{D}$ which, on every message that is $H_{\mathrm{min}}(\mathcal{D})-O(\log(1/\varepsilon))$ bits long, is correct with probability at least $1-\varepsilon$ and is caught with probability at most $\alpha+\delta$ by $\mathcal{W}$.
Finally, we state a known result from information theory in Theorem B.4. This result gives a lower bound on the number of messages that must be sent through independent channels, each of capacity $C$. This is a useful result to recall for the bounds we derive later in Section 5.3.
Theorem B.4 (Capacity limitations).
Let $X$ be a uniform random variable with support of size $q$ (so that $H(X)=\log q$) and $Y^{n}$ be $n$ messages sent through a channel of capacity $C$. Then, in order to transmit $X$ with error at most $\varepsilon$, requires that $n\geq\big(\log q-h(\varepsilon)-\varepsilon\log(q-1)\big)/C.$ Notably, when $Y$ has alphabet size $q^{\prime}$ and is uniformly random with probability $p$, then we obtain
Proof.
$H(X\mid Y^{n})\leq h(\Pr[E])+\Pr[E]\cdot\log(q-1).$
Using the mutual information between $X$ and $Y^{n}$,
Appendix C Algorithms
@@
C.1 Inverse Probability Transform (IPT) Fixed-Seed Sampling Likelihood (FSSL) Estimators
@@
Algorithm 3 Token-IPT-DiFR (from model $\theta$ and context $\mathcal{H}$)
@@
1:model parameters $\theta$;token probabilities $p\in\Delta^{|\mathcal{T}|}$; observed token $t^{\star}\in\{1,\dots,|\mathcal{T}|\}$; seed $\mathsf{seed}$; context position $i$; width $\sigma>0$; small $\varepsilon>0$
@@
9:return $S$
Figure 10: Diagram of the Token-IPT-DiFR method. The Token-IPT-DiFR estimates the likelihood of a token under the inverse probability transform by integrating a Gaussian over the token’s probability interval.
C.2 Gumbel-Max FSSLEstimators (FSSL-GM)